[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Sep 15, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Sep 16 02:58:58 UTC 2017


HSU Shift Log for Fri Sep 15 2017 22:58:52 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: The earlier (19z) issue with ATCF communicating with WCOSS  
appears to be resolved. No problems this cycle (03z).

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170916/78a741f7/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list